Circular No. 7700

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 1994ao AND 1995bf A. Gal-Yam, K. Sharon, and D. Maoz, Tel Aviv University, report:

“Following our discovery of SN 1999gv (IAUC 7405) in deep, duplicate observations of the galaxy cluster Abell 2218 with the Hubble Space Telescope (HST), we report the discovery of two more apparent supernovae in a similar analysis performed on deep HST images of galaxy clusters. We have compared HST images of the galaxy cluster 1604+4304 (z = 0.89) that we obtained during 1995 Feb. 28–Mar. 2 (total exposure time 32 000 s through the F814W filter) with similar images (same exposure time and filter) obtained during 1994 Feb. 25–27. An apparent supernova with magnitude [I814] = 26.5 (equivalent to Cousins I = 26.2), designated SN 1995bf, is detected at α = 16^{h}04^{m}20^{s}.4, δ = +43^{o}04′44′′ (equinox 2000.0). SN 1995bf is absent from the 1994 data (limiting I814 mag 29.5), and it is located 0′′.6 west and 2′′.9 north of an Sb galaxy with z = 0.985 (number 19 of Lubin et al. 1998, A.J. 116, 607). This new source has a stellar point-spread function (PSF) at the HST resolution limit, appearing on 12 individual sub-exposures taken over a time span of 57 hr. We estimate an upper limit of 0′′.025/day on any possible motion of this object. The combined data from 1994 Feb. 25–26 (eight sub-exposures with total exposure time 16 000 s) reveal a source with [I814] = 28.4 (equivalent to Cousins I = 28.1), which is confirmed with the combined data obtained during 1994 Feb. 27 (eight sub-exposures with total exposure time 16 000 s) and is absent from the 1995 combined image (limiting I814 mag 29.5). This apparent supernova, designated SN 1994ao, is detected at α = 16^{h}04^{m}23^{s}.3, δ = +43^{o}03′56′′. SN 1994ao is located 1′′.1 west and 1′′.4 south from an S0 galaxy with z = 0.496 (number 28 of Lubin et al.). This new source has a stellar PSF at the HST resolution limit and shows no motion during three days of observations. We estimate an upper limit of 0′′.05/day on any possible motion of this object. More details and images are available at the following URL: http://wise-

obs.tau.ac.il/~avishay/hst-archive-project/index.html.”

SUPERNOVA 2001dy IN MCG +04-40-16 W. M. Wood-Vasey, UCB and Lawrence Berkeley National Laboratory,

reports that unfiltered CCD images were taken of the field of SN 2001dy (cf. IAUC 7693) with the 1.2-m NEAT/MSSS telescope at Haleakala on July 27.48 UT, but they show no sign of the supernova (limiting mag 19.5).

2001 August 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green